OpenStudy (anonymous):

Seven times a number plus three times another number equals negative one. The sum of the two numbers is negative three. What are the two numbers? Select one: a. 2 and –5 b. –8 and 5 c. –2 and –1 d. –9 and 6

OpenStudy (anonymous):

@FriedRice

OpenStudy (anonymous):

well its saying that 7 times a number plus three times another number = negative one 7 * x + 3 * x = -# 7x + 3x = -#

OpenStudy (anonymous):

so do i have to add 7 to h sides

OpenStudy (anonymous):

and the sum of two numbers = -3 x + y = -3

OpenStudy (anonymous):

then use substitution or elimination method

OpenStudy (anonymous):

7x +3(-x - 3) = -1 7x - 3x - 9 4x - 9 = -1 4x = 8 x = 2

OpenStudy (anonymous):

well then the answer should be A because that the only only answer with a positive 2

OpenStudy (anonymous):

A